OpenStudy (anonymous):

sin 4π 7 cos 5π 21 − cos 4π 7 sin 5π 21

OpenStudy (bradely):

The given expression is in the form sin Acos B-cos AsinB = sin(A-B) A = 4pi/7 and B = 5pi/21 =sin (4pi/7-5pi/21) =sin(pi/3) =sqrt3/2 source: http://www.mathskey.com/question2answer